In a class A of 25 students, 20 passed with 60% or more marks; in another class B of 30 students, 24 passed with 60% or more marks. In which class was a greater fraction of students getting with 60% or more marks?

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Total number of students in Class A = 25 Students passed in first class in Class A = 20
Hence, fraction = 20 / 25
= 4 / 5
Total number of students in Class B = 30 Students passed in first class in Class B = 24
Hence, fraction = 24 / 30
= 4 / 5
∴An equal fraction of students passed in first class in both the classes

​Thanks
Shraddha Singh
2 years ago
Dear Student,
 
No. of students in class A = 25
Students passed in class A = 20
So, fraction = 20/25 = 4/5
 
No. Of students in class B = 30
Students passed in class B = 24
So, fraction = 24/30 = 4/5
 
Therefore, an equal fraction of students passed in both the classes.
 
Hope it helps!
Thank you